OpenStudy (anonymous):

If we perform the experiment "Toss two dice", what is the probability of getting a sum of two? A.) 1/6 B.) 1/12 C.) 1/36

OpenStudy (anonymous):

possible outcomes = 6*6 =36 favorable outcomes = (1,1) = 1 probability = no of favorable outcomes/total no of outcomes = 1/36